Show that acceleration is directly proportional to the displacement in simple pendulum
1
Expert's answer
2020-10-13T09:35:43-0400

Let the equation of SHM, 


"x=A\sin(ωt)\\v=Aω\cos(ωt)\\a=−Aω ^
2
 \sin(ωt) ⇒a=−ω^ 
2
 x"

Acceleration is directly proportional to displacement of the particle from the mean position.